OpenStudy (anonymous):

Biophysics, population kinetics... I have trouble understanding what my prof used to derive the Laplacian regarding population equations. Can someone here help me figure it out? (The equations follow in next post)

OpenStudy (anonymous):

The course uses the example of lynxes-hare connection to explain the equations. First the initial part I get: \[(1)\frac{ dN _{H} }{dt }= k _{H}N _{H}-k _{HL}N _{L}N _{H}\] \[(2)\frac{ dN _{L} }{dt }= k _{LH}N _{H}N _{L}-k _{L,dth}N _{L}\] stationary equilibrium (these I figured out myself) \[\frac{ dN _{H} }{dt }=0= k _{H}N _{H}-k _{HL}N _{L}N _{H} -> N _{L}=\frac{ k _{H} }{ k _{HL} }\] \[\frac{ dN _{L} }{dt }= 0=k _{LH}N _{H}N _{L}-k _{L,dth}N _{L}-> N _{H}=\frac{ k _{L,dth} }{k _{LH} }\] N_h: number of hares N_lL: number of lynxes k_h. N_h: number of births for hares k_hl . N_l. N_h: number of hares killed by lynxes The course continues to including the hunt on hares and gives the following equations \[(3) N _{L,dth}=\frac{ (k _{H}-k _{H,dth}) }{ k _{HL} }\] and \[N _{H,dth}=\frac{ k_{L,dth} }{ k_{LH} }\] So it seems to me that the stationary equations I found are the death equations, except that the hare's also includes human hunting, and not just lynxes hunting them. Now the part that confuses me: we investigate what happens in case of small distortions to the equilibrium \[\Delta N _{H}= N _{H}-N _{H,dth}\] \[\Delta N _{L}= N _{L}-N _{L,dth}\] Then my course says "from (1) (2) and (3) it follows that" \[(4) \frac{ d (\Delta N _{H})}{ dt}=-\frac{ k _{HL} k _{L,dth}}{ k _{LH} }\Delta N _{L}\] \[(5) \frac{ d (\Delta N _{L})}{ dt}=\frac{ k _{LH} k _{H}}{ k _{HL} }\Delta N _{H}\] The course does mention that for the last two she uses... \[\frac{ dN _{H} }{ dt }=f _{1}(N _{H},N _{L})\] \[\frac{ dN _{L} }{ dt }=f _{2}(N _{H},N _{L})\] and so \[\frac{ d (\Delta N _{H})}{ dt }=\left( \frac{ \delta f _{1} }{ \delta N _{H} } \right)_{dth} \Delta N _{H}+\left( \frac{ \delta f _{1} }{ \delta N _{L} } \right)_{dth} \Delta N _{L}\] \[\frac{ d (\Delta N _{L})}{ dt }=\left( \frac{ \delta f _{2} }{ \delta N _{H} } \right)_{dth} \Delta N _{H}+\left( \frac{ \delta f _{2} }{ \delta N _{L} } \right)_{dth} \Delta N _{L}\] But when I'm confused which equations she uses exactly, because when I use (1) and (2) I don't end up with (4) and (5).